Tech Staff Meeting Notes June 6, 2016

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


News & Highlights

Jim and Dave will be attending NetOpsVIII in Pasadena on October 11-12. More info here:

James is making progress on SIS. After initial load of dataless, need to schedule a training session. John will assist loading hardware for ET.


2016 Preventive Maintenance 0% complete

We have an obsidian to replace a CMG5TD. Dave will select the most convenient one for upgrade. Do they use the same rock2slink/slink2ew data transport?


Data taken from:


Misc Problems
CPAR- spikes
EPRM - dropping out
HOVM -spikes and bad EHN see image
HDAR2.*.20 - see image
FCAR.HNE - see image
SEAR - see image

MORT is closed. Need to look for a compromise site to fill the hole left by it and HAWT. Needs to assume flooding. Not a top priority at this time.

Have five new ISIS3 upgrades ready to go that will complete the NMAD node. Three more are on order.

Granite HD2 from the I40 Bridge is in house with a bad flash card. Steve is getting an industrial grade replacement with the appropriate Kinemtrics image. He'll obtain several additional industrical grade cards to flash the image too. The recorder will be replaced when we do the instrument replace later this summer.

There was an email discussion on the history of OXF which led to digging up the MOU from 1993. Mitch checked with Dave Wilson at ASL and it may be possible to replace the MOU, which is endorsed at a fairly high level within USGS, with a "station agreement" which is less formal.

NODE Tasks

All nodes are due a firmware upgrade that will sufficiently change things, the the initial upgrade and configure should be done at CERI, then swap hardware at the node.

John checked other networks one plans to replace Windows digiziting boxes and found none. NC has migrated to the NI X-series boards and adsend does work with it though its still on MS Windows. John will try to find out which specific and OS is used.

DBB: Move the DBB link at CERI from the west end of 3892 to the east side of 3904 and eliminate the thin-wire. Tower is up to test the waters.



The generator at LNXT (now officially known as The Money Pit) has a bad transfer switch. Dave will obtain a new one ($3100) and Jim et al. will install. Then need to get a WASL or KLUV radio engineer to participate in a load transfer test.

Uverse is likely avaialable at the lnxt node. Steve will investigate to see if its cost effective to upgrade from dsl. We don't need the band width but service for dsl will continue to deteriorate. Current costs for lnxt dsl is $119.55/month.




Backup Liebert has a bad blower. Workorder submitted

Flooding in the expansion space has become a problem. A Workorder was submitted and P-plant engineers were on-site to assess the problem. The most likely solution will be a retention wall and removal of material to make the foundation above grave.


Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Noise study picked up compressor noise from a nearby air quality station. Need to add isolation mounts and test farther away. We had to return the compact so will need to use a 6TD instead. We need this cell modem back by June. Noise test to begin tomorrow 5/4.




Misc Problems

GMG has had a power upgrade, new batteries, new solar contoller, and better surge suppression. And it still went down on 1/18. James can get to the radio at the station which is on a separate power feed than the reftek. Jim will send a new BC10 and John will investigate.

John's updated todo list .


Microwave has degraded bandwidth due to increased tree height. Can move receive from EMA to Etowah Utility building. Alternatively Dalton Communications is building a communications system from Star Mt to the Utility building and they are willing to give us a static ip and use their comcast business interenet to get out of Etoway if they have sufficient bandwidth. John is looking into it.

Improve RCGA AC power and perform maintenance work on shed.

Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are t oo tall to get a signal out. Looking for an alternate site.


HPNC node is up. Need to fix Win7 to allow earthworm commands (e.g. status, restart etc) from the ssh command line. 9 channels(GFM, MGNC, and DSNC) are being exported over a cell modem temporarily until internet is established. Still waiting on completion of communcations to get on local internet. We need this modem back before June 30. John has a Thursday 5/5 meeting to do this.

Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.

GFM - Temp station has a bad battery with very sporadic data availability. Horizontals are noisy possibly because of leveling issues. John will replace the battery and we'll then be able to better test for noise and transients.


Replace VHTN E-W and N-S Instruments.

WMTN pit needs to be moved due to erosion.

We're still repeating CPRT through GTTN. The new internet host at the Corryton Senior Center for CPRT has been formally approved by Greg Hill the center coordinator and the Knox County facility supervisor.


Bob has funding for a new CUS station. Deadline is Aug 31 for obligation of all funds. It will be a drill braced monument rather than the piling style. MODOT already has a good quality station near Hayti so Chris is now looking at Ridgely. Have a perfect site with a landowner who already hosts seismic stations but its been difficult getting in contact with him. Chris is working on it. Need a lat/lon so that we can start the bidding.

CVMS has intermittent connection problems with wireless. Its hardwired for now but the wireless needs to be replaced to preserve electical isolation between the building and vault.

HCES got an antenna swap but suffers from the usual outages after school lets out for the summer.

AG Network

TA/AO adoptions. U40A and X40A will become AG stations. Have refteks in house to replace the Q330s. Have borrowed cell modems from Scott we can use until we liberate the modems from ET. If AGS obtains funds for O&M we'll use AGS modems and keep only one CERI modem for a floater. Takeover is scheduled for the week of June 20.


Needs some clean up.

The FDL wiki is here:


The seisnet group on umdrive is at You will need your uuid and password.

The seisnet wiki is at You will need your uuid and password.